OpenStudy (anonymous):

Good afternoon. :) Ok i need help with this problem: Solve for P: 20p+5p-20=21p+4 PLEASE help. :D (Please include how you answered problem. I've been having trouble with it so I want to see)

OpenStudy (nubeer):

25p -21 p = 20+4 4p = 24 p = 6

OpenStudy (anonymous):

^ THANKSS! :D

OpenStudy (nubeer):

welcome

OpenStudy (anonymous):

:D

OpenStudy (nubeer):

:)